Circular No. 6492

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bl IN ANONYMOUS GALAXY C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discov-

ery, on two films taken on Oct. 11.9 and 12.9 UT with the OCA Schmidt telescope, of a supernova (V ∼ 17.4) located at α = 0^{h}36^{m}17^{s}.97, δ = +11^{o}23′40′′.5 (equinox 2000.0), which is 3′′.0 west and 5′′.8 north of the galaxy’s center. Nothing is visible at this position on the original Palomar Sky Survey or on a IIIa-J plate taken in 1987. A nearby star (V ∼ 16.0) is at position end figures 19^{s}.29, 24′09′′.4.

P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that spectra obtained by P. Berlind on Oct. 16.3 UT with the 1.5-m Tillinghast telescope show that this is a type-Ia supernova at maximum light. Hα emission from the host galaxy indicates a redshift of z = 0.036. CCD images taken with the Whipple Observatory 1.2-m reflector by L. Macri show SN 1996bl to be 2′′.4 west and 5′′.0 north of the galaxy’s center at magnitude V = 17.0.

SUPERNOVA 1996bm IN ANONYMOUS GALAXY The Mount Stromlo Abell Cluster Supernova Search Team (cf. IAUC

6418) reports the discovery of an apparent supernova on a CCD image taken by R. Soria on the Mount Stromlo 1.27-m telescope (+ Macho Camera). SN 1996bm is located near Abell 2819 at α = 0^{h}44^{m}40^{s}.85, δ = −63^{o}44′44′′.6 (equinox 2000.0), which is 0′′.3 west and 0′′.6 north of the nucleus of the host galaxy. SN 1996bm is visible at mag ∼ 20.5 on V and R CCD images taken on Oct. 4, and was not detected on similar frames taken on Sept. 1; a nearby HST guide star of mag 14.1 has position end figures 39^{s}.35, 42′16′′.7. CCD images taken with the 1.7-m reflector on Oct. 10 show the supernova to be marginally brighter than on the discovery frame. Finder charts can

be found at http://msowww.anu.edu.au/~reiss/Abell SNSearch.

1996 SK D. Rabinowitz, Jet Propulsion Laboratory, writes: “Observations of

this Apollo-type object (cf. MPEC 1996-T02), obtained on Oct. 15 and 16 UT with the 0.61-m reflector at Table Mountain, yield V − R ∼ +0.73 ± 0.04 and V − I ∼ +0.99 ± 0.04. Assuming V − R = +0.367 and V − I = +0.705 for the sun, 1996 SK reflectance colors are V − R = +0.40, V − I = +0.29. This is much redder than any main-belt asteroid. Follow-up observations on a larger telescope are urged.”

1996 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green